OpenStudy (anonymous):

Simplify. Assume the divisor is not zero. (8x2y2 – 4xy) ÷ 4xy A. 2xy – y B. 2xy C. xy – 1 D. 2xy – 1

OpenStudy (anonymous):

Nurali Kinda need help in my history do i type it in the history?

OpenStudy (nurali):

\[\frac{ 8x^2y^2 -4xy}{ 4xy }\] \[\frac{ 8x^2y^2 }{ 4xy }-\frac{ 4xy }{ 4xy }\] \[2x ^{2-1}y ^{2-1}-1\] \[2xy – 1\]

OpenStudy (dwelsher):

2xy-1